OpenStudy (anonymous):

show how to work 10 divided by 3 1/3

OpenStudy (jdoe0001):

keep in mind that \(\large \bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anonymous):

not sure that helps...sorry